# ProgRAG: Hallucination-Resistant Progressive Retrieval and Reasoning over Knowledge Graphs

### v1

@@ Introduction
Figure 1: Comparison between existing KG-enhanced LLMs and the proposed framework.
Recent advancements have led to the development of KG-enhanced LLMs (Pan et al. 2024), which integrate KGs in either the fine-tuning or inference phases. Despite these improvements, LLM fine-tuning methods (Mavromatis and Karypis 2024; Luo et al. 2023b, a) often require substantial computational resources and struggle to generalize effectively to unseen knowledge. A complementary line of methods using KGs only in the inference phase, i.e., KG-enhanced LLM inference methods, can be broadly categorized into two approaches in terms of retrieval, as illustrated in Figure 1(a): (1) LLM-as-retriever approach in which the LLM itself guides the retrieval process by leveraging its internal knowledge, and (2) external retriever-based approach in which an external retriever extracts relevant reasoning paths or subgraphs. Each approach relies on different sources of knowledge, i.e., internal or external, to retrieve supporting evidence for reasoning.
@@
For the sub-question $q_{i}$, let $T_{re}(q_{i})$ denote the set of triples remaining after top-$p$ sampling. From $T_{re}(q_{i})$, we aim to obtain answer triples, with the source entity $s_{i}$ as the head and the predicted answer to $q_{i}$ as the tail. Specifically, all triples $t\in T_{re}(q_{i})$ are sorted in descending order of $u_{t}$ and provided to the LLM along with $q_{i}$. [^3] If the LLM identifies plausible answers in $T_{re}(q_{i})$, it then outputs them; otherwise, it returns “None”, indicating insufficient evidence. In such cases, ProgRAG returns to the relation pruning step with the unused relations in $R_{re}(q_{i})\setminus R_{pr}(q_{i})$ to explore alternative paths, thereby enabling early termination of incorrect reasoning and reducing redundant inference. ProgRAG then assesses the reliability of the LLM outputs to detect its potential hallucinations, preventing error propagation in multi-hop reasoning, which is called “Uncertainty Quantification”, i.e., ProgRAG measures Aleatoric Uncertainty (AU) of the LLM outputs based on evidential modeling (Sensoy, Kaplan, and Kandemir 2018; Ma et al. 2025). At the generation step $c$ in which the LLM outputs the first token of answer entities, the top-$K$ logits are selected and used as parameters of a Dirichlet distribution:
$\alpha_{k}=\text{LLM}(\tau_{k}|q_{i},\boldsymbol{a}_{c-1}),\alpha_{0}=\sum_{k=1}^{K}\alpha_{k}$ (4)
$\text{AU}(a_{c})=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}(\Psi(\alpha_{k}+1)-\Psi(\alpha_{0}+1))$ (5)
where $\boldsymbol{a}_{c-1}$ is the token sequence generated up to step c1c-1, $\alpha_{k}$ ($1\leq k\leq K$) represents the top-$K$ logits, $\tau_{k}$ is the token with the $k$-th highest logit, $\alpha_{0}$ is the total evidence of the Dirichlet distribution, and $\Psi$ is the digamma function. The resulting $\text{AU}(a_{c})$ serves as the uncertainty estimate for the LLM response. If it exceeds a predefined threshold, indicating low model confidence, the response is refined with the top-$l$ triples from $T_{re}(q_{i})$ as external evidence. Otherwise, the response is accepted as is. If $q_{i}$ is the final sub-question, the model proceeds to the subsequent stage; otherwise, it continues to the next sub-question $q_{i+1}$.
@@ Appendix I Illustration of Question Decomposition
Our implementation of ProgRAG is available at https://anonymous.4open.science/r/ProgRAG-F4C2. The code is structured as follows:

### v2

@@ Introduction
Figure 1: Comparison between existing KG-enhanced LLMs and the proposed framework. E, I, and C denote External, Internal, and Core Supporting Evidences, respectively, and UQ in (b) indicates uncertainty quantification.
Recent advancements have led to the development of KG-enhanced LLMs (Pan et al. 2024), which integrate KGs in either the fine-tuning or inference phases. Despite these improvements, LLM fine-tuning methods (Mavromatis and Karypis 2024; Luo et al. 2023b, a) often require substantial computational resources and struggle to generalize effectively to unseen knowledge. A complementary line of methods using KGs only in the inference phase, i.e., KG-enhanced LLM inference methods, can be broadly categorized into two approaches in terms of retrieval, as illustrated in Figure 1(a): (1) LLM-as-retriever approach in which the LLM itself guides the retrieval process by leveraging its internal knowledge (corresponding to region I in the figure), and (2) external retriever-based approach in which an external retriever extracts relevant reasoning paths or subgraphs (corresponding to region E). Each approach relies on different sources of knowledge, i.e., internal or external, to retrieve supporting evidence for reasoning.
@@
For the sub-question $q_{i}$, let $T_{re}(q_{i})$ denote the set of triples remaining after top-$p$ sampling. From $T_{re}(q_{i})$, we aim to obtain answer triples, with the source entity $s_{i}$ as the head and the predicted answer to $q_{i}$ as the tail. Specifically, all triples $t\in T_{re}(q_{i})$ are sorted in descending order of $u_{t}$ and provided to the LLM along with $q_{i}$. [^3] If the LLM identifies plausible answers in $T_{re}(q_{i})$, it then outputs them; otherwise, it returns “None”, indicating insufficient evidence. In such cases, ProgRAG returns to the relation pruning step with the unused relations in $R_{re}(q_{i})\setminus R_{pr}(q_{i})$ to explore alternative paths, thereby enabling early termination of incorrect reasoning and reducing redundant inference. ProgRAG then assesses the reliability of the LLM outputs to detect its potential hallucinations, preventing error propagation in multi-hop reasoning, which is called “Uncertainty Quantification”.
ProgRAG measures the aleatoric uncertainty (AU) of the LLM outputs by applying evidential modeling (Sensoy, Kaplan, and Kandemir 2018; Ma et al. 2025) to the top-$K$ logits at the generation step in which the LLM produces the first token of the answer entity. If the uncertainty exceeds a predefined threshold, indicating low model confidence, the response is refined with the top-$l$ triples from $T_{re}(q_{i})$ as external evidence. Otherwise, the response is accepted as is. If $q_{i}$ is the final sub-question, the model proceeds to the subsequent stage; otherwise, it continues to the next sub-question $q_{i+1}$.
@@ Appendix I Illustration of Question Decomposition
The code is structured as follows: